OpenStudy (anonymous):

Calculate the value of P, such that the line x-2/6=y-1/P=z-5/4 is perpendicular to the plane 3x-y-2z=0

OpenStudy (anonymous):

Do you mean \[\frac{x-2}{6}=\frac{y-1}{P}=\frac{z-5}{4}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

There is one clumsy way. The equation for the line implies there exists a parametric equation in the form \[\vec{r}(t)=t(4,P-1,-1)+(2,1,5)\] and the equation of the surface satisfies the equation\[\vec{X}(t,s)=(t,s,\frac{3t-s}{2})\]Find the direction of the normal of the surface by\[\vec{n}=\frac{\partial \vec{X}}{\partial t}\times\frac{\partial\vec{X}}{\partial s}\]This means \[(4,P-1,-1)=\lambda\vec{n},\lambda\in R\]Find λ and solve for P. This is definitely not the best way to do it, maybe someone else will give you a simpler version. :P

OpenStudy (anonymous):

let the point on the line be... \[x=6t+2\quad y=Pt+\quad z=4t+5\] eq of normal to the plane: \[n⃗ =<3,−1,−2>\] this is the direction of the given line.

OpenStudy (anonymous):

Upon some Google-ling, for plane with equation ax+by+cz+d=0 passing through (x1,y1,z1), i.e. (0,0,0) in this case, have the equation (x-x1)/a=(y-y1)/b=(z-z1)/c. Then we have b=-1=P

OpenStudy (anonymous):

do we know that the given line passes throught (0,0,0)?

OpenStudy (anonymous):

We know the plane passes through (0,0,0)

OpenStudy (anonymous):

@electrokid what to do next

OpenStudy (anonymous):

@msingh It is done already, P=-1

OpenStudy (anonymous):

@agostino oh yea.. duh

OpenStudy (anonymous):

@electrokid lol

OpenStudy (anonymous):

i didn't get it from where @electrokid has left

OpenStudy (anonymous):

@msingh forget me.. use agostino's method mentioned in the most recent post

OpenStudy (anonymous):

after this step, plse explain let the point on the line be... x=6t+2y=Pt+z=4t+5 eq of normal to the plane: n⃗ =<3,−1,−2> this is the direction of the given line.

OpenStudy (anonymous):

i m not able to understand , plse anyone explain

OpenStudy (amistre64):

what do you know about planes?

OpenStudy (anonymous):

it is infinte set of poinnts

OpenStudy (anonymous):

forming a flat surface which is extended infinitely far in alll directions

OpenStudy (amistre64):

good, this flat surface can be defined by a vector that is perpendicular to all the vectors in the plane. This perp vector is called a normal.

OpenStudy (amistre64):



OpenStudy (amistre64):

there is a property of vectors such that there dot product equals zero if the vectors are per to each other.

OpenStudy (amistre64):

lets define the a vector in the plane as a directed line from point(x,y,z) to point(xo,yo,zo) all the vectors then take the form <x-xo, y-yo, z-zo> the vector that is perp to all these vectors in the plane is the normal <a,b,c> the dot product of the plane and normal vectors is then a(x-xo) + b(y-yo) + c(z-zo) = 0